Which of the following statements describes the following expression if the sum is compounded annually for 2 years?
A = 1350(1.06)2

Answer by jim_thompson5910(35256)

The rest of your problem is cut off, but I'm assuming your teacher is aiming for something like "The amount $1,350 is deposited into an account with 6% interest. The money is compounded annually for 2 years".

Note: this is assuming no other deposits or withdrawals take place. Also, this assumes the interest rate stays at 6%. Keep in mind that 6% = 6/100 = 0.06. To get 6% growth, we'd have 1+r = 1+0.06 = 1.06